College of the Holy Cross, Fall Semester, 2011

MONT 102N, Final Exam Wednesday, December 14, 8:00 AM

Professor Little

Your Name:

Information: Do all work on these sheets. You may use the back of a sheet if necessary, but please alert me with a note if you do. In a case where a later part of a question depends on a previous part, I will give credit for the later part as long as you have a correct approach, even without the correct value from the previous part.

Please do not write in the space below

Problem Points/Poss
I  / 40
II  / 45
III  / 25
IV  / 30
V  / 60
Total  /200

Have a safe and joyous holiday season!